OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. A) y = -3x - 13 B) y = -3x + 11 C) y = -3x + 13 D) y = -3x + 1

OpenStudy (anonymous):

Use point-slope form \[y - y_1 = m(x - x_1)\]\[y + 5 = -3(x - 2)\]\[y = -3(x - 2) - 5\]Just simplify now.

OpenStudy (anonymous):

y=mx+c where m is your slope. so all of the answers have the right slope. just sub the point (2,-5) in so in D) -5=(-3)(2)+1 which gives -5=-5 so your answer is D) :)

OpenStudy (anonymous):

ohhh ok thanks so much your a life saver(: math is not my strong point what so ever(:!

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

but actually do you think you could help me with another question im having trouble with? please